# DiaHalu: A Dialogue-level Hallucination Evaluation Benchmark for Large Language Models

### v1

@@ 4.3 Human Annotation
Annotating the hallucination and its types in the benchmark is a very challenging task. Since there may be more than one instance of hallucination in multi-turn dialogues. Also, some hallucination types in edge cases are difficult to differentiate. Therefore, the entire annotation process demands a high level of expertise from annotators and requires detailed definitions for ambiguous contents.
@@
of our benchmark are all seasoned researchers in the field of linguistics and natural language processing. The annotators all obtain at least a bachelor’s degree, pass the IELTS or TOEFL exams, and are proficient in using search engines such as Google and Bing. In addition, everyone exhibits strong collaborative and communicative skills. To ensure the validity of the annotation information, we invite experienced experts in the field of LLMs’ hallucination detection from both academia and industry to conduct random sampling checks.
@@
is divided into three steps. (1) Each annotator labels 50 samples for each domain, followed by a careful discussion. The annotators are required to label the presence of hallucination, hallucination types and locations, along with the corresponding explanations. The discussions intricately define the application scope of each hallucination label (Appendix A.4), and ultimately we decide to annotate the corresponding hallucination type on the first dialogue round in which hallucination occurs. (2) All the annotators label the entire dataset, discussions and corrections are made for inconsistent annotations. (3) Statistical analysis is performed on the annotated results. For more details, please refer to Appendix A.5.
@@
For evaluating the inter-annotator consistency, we calculate the Fleiss’s Kappa (Randolph, 2005) of Inter-Annotator Agreement (IAA) (Artstein, 2017), which is a statistical measure used to assess the degree of agreement among multiple raters for a set of items.
@@
According to the above formulas, the calculated result for Fleiss’s Kappa of our benchmark is 0.8709, representing almost perfect agreement among all the annotators.
Annotation Results
After annotating the entire dataset, we conduct several statistical analyses on it. Table 2 reveals the probability of hallucination occurring in each dialogue domain. The results indicate that hallucinations are highly likely to arise in Knowledge-grounded dialogue and reasoning dialogue. Therefore, the knowledge and reasoning abilities of LLMs still need further improvement. Despite LLMs’ powerful multi-turn dialogue capability, faithfulness hallucination such as irrelevance, incoherence and overreliance still persists. Figure 5 presents the proportion of each hallucination type in each dialogue domain. Irrelevance, incoherence, and overreliance Widely exit in daily dialogue contexts, such as task-oriented and chit-chat scenarios. In knowledge-grounded dialogue, the factuality hallucination constitutes a significant proportion, while in reasoning dialogue, almost all hallucinations are defined as errors in reasoning. This statistical information can help us understand the types of hallucination in LLMs’s multi-turn dialogues, facilitating an exploration of their origins and contributing to the elimination of these types of hallucination.
Knowledge Task Chit Reasoning Overall
# Number
272
131
186
159
748
# Non-Halu
141
84
115
79
419
# Halu
131
47
71
80
329
Halu Rate (%)
48.16
35.88
38.17
50.31
43.98
Table 2: The statistical information of hallucination on the four dialogue domains.
Figure 5: The distribution of five different hallucination types within the four dialogue domains.
Random 48.19 30.53 37.38 35.71 21.28 26.67 32.50 18.31 23.42 50.00 23.75 32.20 43.39 24.92 31.66
$SelfCheckGPT_{B}$ 47.92 17.56 25.70 32.14 19.15 24.00 37.93 15.49 22.00 57.50 28.75 38.33 45.52 20.06 27.85
$SelfCheckGPT_{N}$ 64.41 29.01 40.00 34.25 53.19 41.67 46.58 47.89 47.22 64.00 20.00 30.48 49.13 34.35 40.43
$SelfCheckGPT_{P}$ 56.45 26.72 36.27 44.73 36.17 40.00 43.33 54.92 48.45 60.96 62.50 61.73 51,84 42.86 46.92
LLaMA-30B 38.46 3.82 6.94 40.00 4.26 7.69 50.00 11.27 18.39 71.43 6.25 11.49 48.78 6.08 10.81
Vicuna-33B 50.00 4.58 8.39 100.00 4.26 8.16 28.57 2.82 5.13 41.38 15.00 22.02 44.00 6.69 11.61
ChatGPT3.5 50.00 0.76 1.50 0.00 0.00 0.00 42.86 4.23 7.69 55.56 6.25 11.24 45.00 2.74 5.16
ChatGPT4 82.46 35.88 50.00 73.68 29.79 42.42 56.52 18.31 27.66 71.43 81.25 76.02 73.16 42.25 53.56
Table 3: The classification results on four kinds of baselines, and the best F1 scores are in bold form.
ChatGPT3.5 50.00 0.76 1.50 0.00 0.00 0.00 42.86 4.23 7.69 55.56 6.25 11.24 45.00 2.74 5.16
w/ CoT 45.45 3.82 7.04 20.00 2.13 3.85 28.57 2.82 5.13 53.85 8.75 15.05 41.67 4.56 8.22
w/ retrieval 70.00 5.34 9.93 - - - - - - 71.43 12.50 21.27 - - -
ChatGPT4 82.46 35.88 50.00 73.68 29.79 42.42 56.52 18.31 27.66 71.43 81.25 76.02 73.16 42.25 53.56
w/ CoT 82.02 55.73 66.36 72.41 44.68 55.26 69.05 40.85 51.33 69.52 91.25 78.92 73.96 59.57 65.99
w/ retrieval 75.90 48.09 58.88 - - - - - - 70.65 81.25 75.58 - - -
Table 4: The results of CoT and retrieval techniques which can enhance hallucination detection on ChatGPT3.5 and ChatGPT4.
5 Experiments
In this section, we assess the performance of several evaluation models and methods on the dataset we introduced. Thereby, we can determine the effectiveness of existing methods in detecting dialogue-level hallucination.
5.1 Baselines
We select some powerful LLMs by providing specific prompts, to assess the presence of hallucinations. Similarly, we also conduct experiments using existing hallucination detection methods.
Random
SelfCheckGPT
LLaMA-30B && Vicuna-33B
They are two well-pretrained and widely deployed open-source LLM backbones [^2] https://huggingface.co/huggyllama/llama-30b, https://huggingface.co/lmsys/vicuna-33b-v1.3 (Touvron et al., 2023; Chiang et al., 2023). We provide a specially designed prompt to assist with detection. More details are shown in Appendix A.6.
ChatGPT3.5 && ChatGPT4
Both of these models are developed by OpenAI [^3] https://chat.openai.com/. ChatGPT3.5 marks the beginning of the era of large language models and GPT4 is currently the most powerful language model (Wu et al., 2023b; OpenAI, 2023). We employ the same prompt as LLaMA and Vicuna do.
5.2 Metrics
Despite we annotating the types of hallucinations in the dataset, achieving consistent labels even among humans requires further discussion. Therefore, similar to past hallucination detection efforts, we merely focus on a binary classification task of determining the existence of hallucination. Consequently, we utilize binary classification evaluation metrics. The positive label for this classification task is set as "Halu", for our main focus is testing the model’s ability to recognize hallucination.
5.3 Main Results
From the results in Table 3, we can get the following conclusions.
First, DiaHalu is a very challenging benchmark for dialogue-level hallucination detection.
Except for ChatGPT4, the F1 scores of all other detection methods do not exceed 50.00. Existing language models, such as LLaMA-30B and Vicuna-33B, are not effective in accurately discerning most samples that involve hallucinations. Regarding the specialized detection method SelfCheckGPT, it achieves relatively better performance when applying the prompt and NLI methods. However, distinguishing between positive and negative samples proves challenging with BERTScore.
Second, when performing self-detection using ChatGPT3.5, there is a noticeable phenomenon of overconfidence.
Our dataset is primarily generated by ChatGPT3.5, which exhibits high confidence in its output. Despite providing a specially designed detection prompt, it still struggles to differentiate whether the dialogue content is hallucinating or not. So, the majority of its output labels are "Non-Halu".
Apart from the specialized hallucination detection method SelfCheckGPT, the results from directly harnessing LLMs for judgment indicate that the recognition accuracy for task-oriented and chit-chat domains of dialogue are much lower than that for the knowledge-grounded and reasoning dialogue. This is because the hallucination types in the knowledge-grounded and reasoning dialogue are primarily Non-factual and Reasoning Error, which present in a more direct and apparent manner. While task-oriented and chit-chat domains mainly consist of three types of faithfulness hallucination, which requires a LLM to possess long-term memory and the ability to recognize topic transitions in dialogues.
5.4 Chain-of-Thought and Retrieval for Detection
Chain-of-Thought and Retrieval are two important techniques for enhancing the reasoning ability of LLMs. Chain-of-Thought (CoT) (Feng et al., 2023a) describes the organized sequence of logical reasoning that unfolds during thinking. Retrieval (Gao et al., 2023) means retrieving relevant contents from the media to supplement external knowledge for LLMs. In this section, we test whether these two techniques can improve the effectiveness of hallucination detection in Table 4. More details are depicted in Appendix A.7.
The experimental results indicate that both methods have facilitating effects on hallucination detection. However, the combination of GPT4 with retrieval methods shows a decrease of around 0.5 points in F1 score in the domain of reasoning detection. We believe that GPT4 inherently possesses a strong reasoning ability, and therefore, additional retrieval contents introduce noise to its judgments.
6 Conclusion
In this paper, we propose a dialogue-level hallucination evaluation benchmark. We construct the benchmark in a three-step process and have the experts annotate all the data. DiaHalu covers four multi-turn dialogue domains and five hallucination subtypes. Experiments through some well-known LLMs and detection methods on the dataset show that DiaHalu is a challenging task, holding significant value for further research (Appendix A.8).
Limitations
This paper proposes a novel dialogue-level hallucination detection evaluation benchmark named DiaHalu. The benchmark covers four multi-turn dialogue domains and five hallucination subtypes. There is significant value for further research. However, two main limitations also exist. (1) During the second step of the benchmark construction phase, aligning the contents of speaker A with human language consumes a considerable amount of time and effort. Frequent calls to the ChatGPT3.5 API_Keys result in a significant expense. Simultaneously, achieving consistency among all annotators led to prolonged discussion time. (2) We do not partition the dataset into training, validation, and test sets. This is because we consider the black-box detection scenario for those closed-source LLMs. If a division into these three types of datasets is required, it would necessitate more data samples and larger resource consumption.
Ethics Statement
The benchmark is primarily generated by ChatGPT3.5. We obtain all the API_Keys through a paid subscription. All the annotators are real and they receive corresponding compensation and rewards. The entire process and outcomes are free from intellectual property and ethical legal disputes.
Acknowledgements
Appendix A Appendices
A.1 System Prompts for Dialogue Generation
In this section, we present the specific form of the system prompts for the four dialogue domains. We use ChatGPT3.5 (gpt-3.5-turbo-1106) [^4] https://platform.openai.com/docs/guides/text-generation/chat-completions-api with the temperature 0.1. The brief system prompts of the knowledge-grounded dialogue, task-oriented dialogue, chit-chat dialogue and reasoning dialogue are respectively presented in Figure 6, Figure 7, Figure 8 and Figure 9.
Figure 6: The brief system prompts for knowledge-grounded dialogue.
Figure 7: The brief system prompts for task-oriented dialogue.
Figure 8: The brief system prompts for chit-chat dialogue.
Figure 9: The brief system prompts for reasoning dialogue.
A.2 Data Format
Figure 10: One sample from the benchmark
A.3 The Comparison with Other Benchmarks
In Table 5, we present the differences between DiaHalu and other hallucination detection benchmarks, highlighting the distinctive features that our DiaHalu is a naturally generated dialogue-level benchmark by LLMs, with various kinds of hallucination and explanations. All the compared benchmarks can be referred to in Section 2.1. HaluEval, WikiBio+, and PHD benchmark come from the paper Li et al. (2023), Manakul et al. (2023) and Yang et al. (2023) respectively.
A.4 The Application Scope of Hallucination Labels
Annotating the hallucination and its types in the benchmark is a very challenging task. One of the reasons is that some hallucination types in side cases are defiant to differentiate. Therefore, in the first stage of annotation, we provide detailed definitions for each hallucination type.

### v2

@@ 4.3 Human Annotation
Annotating the hallucination and its types in this dataset is a very challenging task. Since there may be more than one instance of hallucination in multi-turn dialogue. Also, some hallucination subtypes in edge cases are difficult to differentiate. Therefore, the entire annotation process demands a high level of expertise from annotators and requires detailed definitions for ambiguous contents.
@@
of our dataset are all seasoned researchers in the field of linguistics and natural language processing. We invite experienced experts in the field of LLMs’ hallucination detection from both academia and industry to engage in discussion and conduct sampling checks. For more details about the annotators and the experts, refer to Appendix A.6 (The annotators).
@@
is divided into three steps. (1) Each annotator labels some samples for each domain, followed by a careful discussion between the annotators and the experts. The discussions intricately define the application scope of each hallucination label (discussion results in Appendix A.7). (2) All the annotators label the entire dataset, discussions and corrections are made for inconsistent annotations. (3) Statistical analysis is performed on the annotated results. For more details, please refer to Appendix A.6 (The Annotation Process).
@@
For evaluating the inter-annotator consistency, we calculate the Fleiss’s Kappa (Randolph, 2005) of Inter-Annotator Agreement (IAA) (Artstein, 2017), which is a statistical measure used to assess the degree of agreement among multiple raters for a set of items. The final score of Fleiss’s Kappa is 0.8842, representing almost perfect agreement among all the annotators. For more calculation details, please refer to Appendix A.6 (Label Consistency).
Annotation Results
After annotating the entire dataset, we conduct several statistical analyses on it. Table 3 reveals the probability of hallucination occurring in each dialogue domain. The results indicate that hallucination are highly likely to arise in knowledge-grounded dialogue and reasoning dialogue. Therefore, the knowledge and reasoning abilities of LLMs still need further improvement. Despite LLMs’ powerful multi-turn dialogue capability, faithfulness hallucination such as irrelevance, incoherence and overreliance still persists. Figure 4 presents the proportion of each hallucination subtype in each dialogue domain. Irrelevance, incoherence, and overreliance widely exist in daily dialogue contexts, such as task-oriented and chit-chat scenarios. In knowledge-grounded dialogue, the factuality hallucination constitutes a significant proportion, while in reasoning dialogue, almost all hallucination are defined as errors in reasoning. This statistical information can help us understand the subtypes of hallucination in LLMs’ multi-turn dialogue, facilitating an exploration of their origins and contributing to the elimination of these subtypes of hallucination.
Knowledge Task Chit Reasoning Overall
# Number 371 210 263 259 1103
# Non-Halu 199 135 164 129 627
# Halu 172 75 99 130 476
Halu Rate (%) 46.36 35.71 37.64 50.19 43.16
Table 3: The statistical information of hallucination on the four dialogue domains. ‘# number’, ‘# Non-Halu’,‘# Halu’ and ‘Halu Rate’ represent the number of samples, the number of samples without hallucination, the number of samples with hallucination and the proportion of hallucinated samples.
Figure 4: The distribution of five different hallucination subtypes within the four dialogue domains.
Random 41.57 43.02 42.29 31.86 48.00 38.30 38.46 50.51 43.67 49.61 49.23 49.42 40.72 47.06 43.66
$SelfCheckGPT_{B}$ 42.55 23.26 30.08 35.38 30.67 32.86 30.00 18.18 22.64 60.81 34.61 44.12 43.00 26.47 32.77
$SelfCheckGPT_{N}$ 59.46 25.58 35.77 38.84 62.67 47.96 45.19 47.47 46.30 70.58 18.46 29.27 48.65 34.03 40.05
$SelfCheckGPT_{P}$ 55.22 21.51 30.96 48.00 32.00 38.40 45.00 45.45 45.23 62.37 44.62 52.02 52.90 34.45 41.73
FOCUS 46.11 48.26 47.16 34.09 60.00 43.48 36.56 49.49 42.06 50.56 34.62 41.10 41.49 46.64 43.92
LLaMa-30B 37.50 5.23 9.18 30.77 5.33 9.09 50.00 11.11 18.18 81.25 10.00 17.81 49.33 7.78 13.43
Vicuna-33B 45.45 5.81 10.31 42.86 4.00 7.32 36.36 4.04 7.27 51.35 14.62 22.75 46.75 7.56 13.02
Gemini1.5 PRO 80.00 20.93 33.18 60.00 36.00 45.00 70.37 38.38 49.67 73.63 51.54 60.63 71.49 35.29 47.26
ChatGPT3.5 25.00 0.58 1.14 33.33 2.67 4.93 55.56 5.05 9.26 57.14 6.15 11.11 48.48 3.36 6.27
GPT4 80.89 31.98 45.83 74.19 30.67 43.40 67.74 21.21 32.31 74.07 61.54 67.23 75.21 37.61 50.14
Table 4: The classification results on four kinds of baselines, and the best F1 scores are in bold form. The indices $B$, $N$ and $P$ of SelfCheckGPT denote scoring with BERTScore, with NLI and using prompts, respectively.
Prec Rec F1 Prec Rec F1 Prec Rec F1 Prec Rec F1 Prec Rec F1
FaithCritic w/retrieval 28.26 84.54 42.38 - - - - - - 51.63 79.17 62.50 - - -
Gemini1.5 PRO 80.00 20.93 33.18 60.00 36.00 45.00 70.37 38.38 49.67 73.63 51.54 60.63 71.49 35.29 47.26
w/ CoT 81.25 22.67 35.45$\uparrow$ 69.77 40.00 50.85$\uparrow$ 75.00 36.36 48.98$\downarrow$ 72.92 53.85 61.95$\uparrow$ 74.47 36.76 49.23$\uparrow$
w/ retrieval 86.04 21.51 34.42$\uparrow$ - - - - - - 76.70 60.77 67.81$\uparrow$ - - -
ChatGPT3.5 25.00 0.58 1.14 33.33 2.67 4.94 55.56 5.05 9.26 57.14 6.15 11.11 48.48 3.36 6.27
w/ CoT 45.45 2.91 5.46$\uparrow$ 33.33 2.67 4.94 40.00 4.04 7.34$\downarrow$ 47.06 6.15 10.88$\uparrow$ 43.18 3.99 7.31$\uparrow$
w/ retrieval 70.00 4.01 7.69$\uparrow$ - - - - - - 70.58 9.23 16.32$\uparrow$ - - -
GPT4 80.89 31.98 45.83 74.19 30.67 43.40 67.74 21.21 32.31 74.07 61.54 67.23 75.21 37.61 50.14
w/ CoT 86.05 21.51 34.42$\downarrow$ 73.17 40.00 51.72$\uparrow$ 80.56 29.29 42.96$\uparrow$ 71.43 76.92 74.07$\uparrow$ 75.38 41.18 53.26$\uparrow$
w/ retrieval 77.89 43.02 55.43$\uparrow$ - - - - - - 74.40 71.54 72.94$\uparrow$ - - -
Table 5: The results of CoT and retrieval techniques on the three closed-source LLMs. The $\uparrow$ and $\downarrow$ indicate whether CoT and retrieval can promote improvements in F1 score. We also provide the detection results of FaithCritic.
5 Experiments
In this section, we assess the performance of several evaluation models and specialized methods on the dataset we introduced. Thereby, we can trial the effectiveness of existing methods in detecting dialogue-level hallucination. We still conduct more fine-grained detection and explore whether the phenomenon of hallucination snowballing exists.
5.1 Baselines
We select some powerful LLMs to detect hallucination by providing specific prompts. These models include open-source LLMs: LLaMa-30B (Touvron et al., 2023), Vicuna-33B (Chiang et al., 2023), and some closed-source LLMs: Gemini1.5 PRO (Anil et al., 2023), ChatGPT3.5 (Wu et al., 2023c) and GPT4 (OpenAI, 2023). Similarly, we also experiment on specialized existing hallucination detection methods, such as FaithCritic (Dziri et al., 2022a), SelfCheckGPT (Manakul et al., 2023) and FOCUS (Zhang et al., 2023b). For a detailed description of the above baselines, please refer to Appendix A.8 (I. Baselines Selected).
5.2 Metrics
For hallucination detection, we use standard binary classification to determine whether there exists hallucination (Table 4). Meanwhile, we also conduct more fine-grained hallucination-type recognition to judge the specific subtype of hallucination and use micro-F1 score for all hallucination categories (Table 6). Appendix A.8 (II. Metrics Calculation) provides more thorough explanations.
5.3 Main Results
From the results in Table 4, we can get the following conclusions.
First, DiaHalu is a highly challenging benchmark for dialogue-level hallucination detection.
Except for GPT4, the F1 scores of all other detection methods and detecting LLMs do not exceed 50.00. Existing LLMs, such as LLaMa-30B and Vicuna-33B, are not effective in accurately discerning most samples that involve hallucination. Regarding the specialized detection methods FOCUS and SelfCheckGPT (applying prompt and NLI methods), they achieve relatively better performances. However, it proves challenging with BERTScore for SelfCheckGPT.
Second, ChatGPT3.5 shows a noticeable phenomenon of overconfidence.
Our dataset is primarily generated by ChatGPT3.5, which exhibits high confidence in its output. Despite providing a specially designed detection prompt, it still struggles to differentiate whether the dialogue content is hallucinated or not, not along the samples generated by GPT4. So, the majority of its output labels are "Non-Halu".
Apart from the specialized hallucination detection methods, the results from directly harnessing LLMs for judgment indicate that the recognition accuracy for task-oriented and chit-chat domains of dialogue are much lower than that for the knowledge-grounded and reasoning dialogue. This is because the hallucination types in the knowledge-grounded and reasoning dialogue are primarily Non-factual and Reasoning Error, which present in a more direct and apparent manner. Nevertheless, task-oriented and chit-chat domains mainly consist of three subtypes of faithfulness hallucination, which requires a LLM to possess long-term memory and the ability to recognize topics/roles transition in dialogue.
NF Ic Ir Ov RE ALL
Gemini1.5 PRO 18.97 30.49 11.36 4.76 45.41 26.72
ChatGPT3.5 1.16 4.26 0.00 0.00 9.66 3.93
GPT4 29.38 25.00 5.71 4.65 55.66 32.30
Table 6: Fine-grained hallucination-type recognition F1 scores for three LLMs. ‘NF’, ‘Ic’, ‘Ir’, ‘Ov’ and ‘RE’ stand for Non-factual, Incoherence, Irrelevance, Overreliance and Reasoning Error, respectively. ‘ALL’ represents micro-f1 of all hallucination subtypes.
Figure 5: The proportions of the three dialogue round categories. For example, the three values of R7 denote the proportions of ‘these three categories in the 7th round’ within ‘hallucinated samples that have at least seven rounds dialogues’.
5.4 Chain-of-Thought and Retrieval for Detection
Chain-of-Thought (CoT) and Retrieval are two important techniques for enhancing the ability of LLMs. In this section, we test whether these two techniques can improve the effectiveness of hallucination detection in Table 5. More details are depicted in Appendix A.11.
The experimental results indicate that both methods have facilitating effects on hallucination detection. However, Gemini1.5 PRO and ChatGPT3.5 with CoT show a decrease of around 1.00 F1 points in the chit-chat domain. We believe that these two models inherently lack the ability to recognize faithfulness hallucination, and additional CoT contents introduce noise to their judgments.
5.5 Fine-grained Hallucination-type Recognition
Table 6 shows fine-grained hallucination-type recognition results for three open-source LLMs. We can conclude that ChatGPT3.5 fails the recognition of almost all labels. To some extent, Gemini1.5 PRO and GPT4 have the ability to recognize factuality hallucination and reasoning errors, but they have lower F1 scores for the three subtypes of faithfulness hallucination. This result reveals that faithfulness hallucination remains a pressing issue for LLMs.
5.6 Hallucination Snowballing
In this section, we study the hallucination snowballing phenomenon (Zhang et al., 2023c) in our benchmark. Specifically, for each round of dialog (2-10) in all hallucinated samples, we define three categories: I hallucination that appears in previous rounds and also appears in the current round (Halu&Halu), II hallucination that appears in previous rounds but not appear in the current round (Halu&None) and III hallucination that not appear in previous rounds but appears in the current round (None&Halu). We calculate the proportions of these three categories in Figure 5.
First, I is greater than the other two categories (II and III), which means that hallucinated contents are more likely to generate new hallucinated responses. Second, I shows the most obvious increasing trend, indicating that the probability of hallucination increases with the number of dialogue rounds. These two findings validate the hallucination snowballing phenomenon.
6 Conclusion
In this paper, we propose a dialogue-level hallucination evaluation benchmark named DiaHalu. We construct the benchmark in a three-step process. The DiaHalu covers four multi-turn dialogue domains and five hallucination subtypes. Experiments through some well-known LLMs and specialized detection methods on the benchmark show that it is a challenging task, holding significant value for further research (Appendix A.12).
Limitations
This paper proposes a novel dedicated dialogue-level hallucination detection evaluation benchmark named DiaHalu. The benchmark covers four multi-turn dialogue domains and five hallucination subtypes. There is significant value for further research. However, two main limitations also exist. (1) During the second step of the benchmark construction phase, aligning the contents of speaker A with human language consumes a considerable amount of time and effort. Frequent calls to the ChatGPT3.5 or GPT4 API Keys result in a significant expense. Simultaneously, achieving consistency among all annotators led to prolonged discussion time and money cost. (2) We do not partition the dataset into training, validation, and test sets. The primary purpose of evaluation benchmarks is to assess a models’ capabilities. However, if we divide the dataset into the above three categories, this is about assigning capabilities to models. From the perspective of the two objectives, there is a clear difference. Another reason is that we need to consider the black-box detection scenario for those closed-source LLMs. However, if a division into these three types of datasets is necessarily required, it would require more data samples and larger resource consumption.
Ethics Statement
The benchmark is primarily generated by ChatGPT3.5 or GPT4. We obtain all the API Keys through a paid subscription. All the annotators are real people and they receive corresponding compensation and rewards. The entire process and outcomes are free from intellectual property and ethical legal disputes.
Acknowledgements
Appendix A Appendices
A.1 The Comparison with Other Benchmarks
In Table 1, we present the differences between DiaHalu and other hallucination detection benchmarks, highlighting the distinctive features that our DiaHalu is a naturally generated dialogue-level benchmark by LLMs, with various kinds of hallucination and explanations. All the compared benchmarks can be referred to in Section 2.1. HaluEval, WikiBio+, and PHD benchmark come from the paper Li et al. (2023b), Manakul et al. (2023) and Yang et al. (2023) respectively.
A.2 The Four Dialogue Domains
Knowledge-grounded dialogue
Task-oriented dialogue
Chit-Chat dialogue
involves open-ended and non-goal dialogue (Sun et al., 2021). We provide two LLMs with personas and facilitate a dialogue between them. This approach allows for the evaluation of their memory capabilities, conversational coherence, and relevance to the information being discussed.
Reasoning dialogue
centralizes on the logical reasoning and understanding capabilities of LLMs. Following previous works (Chen et al., 2023b; Li et al., 2023a; Buszydlik et al., 2024; Grover et al., 2024; Zheng et al., 2023; Huang and Chang, 2023), we also treat reasoning errors as a kind of hallucination. We have the models discuss mathematical problems to achieve the answers (Kakarla et al., 2024).
A.3 The Distribution of the Sources for Dialogue Topics
The number and proportion of all dialogue topics in DiaHalu across the 10 topic sources are shown in Figure 6.
Figure 6: The distribution of the 10 sources for dialogue topics.
A.4 System Prompts for Dialogue Generation
In this section, we present the specific form of the system prompts for the four dialogue domains. We use ChatGPT3.5 (gpt-3.5-turbo-1106) [^2] https://platform.openai.com/docs/models/gpt-3-5-turbo and GPT4 (gpt-4-1106-preview) [^3] https://openai.com/gpt-4 with the temperature 0.1. The brief system prompts of the knowledge-grounded dialogue, task-oriented dialogue, chit-chat dialogue and reasoning dialogue are respectively presented in Figure 7, Figure 8, Figure 9 and Figure 10.
Previous works prove the LLMs’ ability to follow complex instructions (Manakul et al., 2023; Guan et al., 2023; Chen et al., 2023b; Mündler et al., 2023; Li et al., 2023b; Wang et al., 2023b; Liu et al., 2023; Li et al., 2024b), including some hallucination tasks. Thus, we reference such kinds of prompts, and then we formulate prompts for our benchmark.
Figure 7: The brief system prompts for knowledge-grounded dialogue.
Figure 8: The brief system prompts for task-oriented dialogue.
Figure 9: The brief system prompts for chit-chat dialogue.
Figure 10: The brief system prompts for reasoning dialogue.
A.5 Generation Details and Data Format
We utilize 8 ChatGPT3.5 and 2 GPT4 API Keys, generating approximately 4000 examples in total initially. We conduct a thorough examination of samples to filter the confused formats or contents not aligned with instructions (system prompts). To ensure a balanced number of positive and negative samples, we manually remove samples with highly similar semantics. Finally, the number of samples generated by ChatGPT3.5 and GPT4 are 748 and 355 respectively. The total number is 1103. The reason we use GPT4 to generate data is to ensure its competitiveness for a long time in the future.
Previous well-known hallucination benchmarks (Manakul et al., 2023; Guan et al., 2023; Chen et al., 2023b; Yang et al., 2023) contain 238, 400, 847, and 300 samples respectively, which is at the similar scale as ours. What’s more, our benchmark is at dialogue level, which contains about 7620 rounds (more than 27600 rounds initially) of interactions in total (6.9120 average rounds per dialogue as shown in Table 2). This indicates a larger volume of data compared to previous benchmarks at the sentence and passage levels.
It is also worth noting that, given that we assume the two subjects of the dialogue are A and B, both A and B are set to be either ChatGPT3.5 or GPT4, and it is not possible for one to be ChatGPT3.5 and the other to be GPT4.
Figure 11: One sample from our benchmark DiaHalu
A.6 The Supplementary Details for Annotation
The annotators
all obtain at least a bachelor’s degree, get a high score in IELTS or TOEFL exams, and are proficient in using search engines such as Google and Bing. The annotators are all seasoned researchers in the field of linguistics and natural language processing. In addition, everyone exhibits strong collaborative and communicative skills. We also invite senior experts in the field of hallucination detection from academia and industry to engage in discussions and data checking.
The Annotation Process
To ensure the annotation quality, we perform three steps for annotation as described in Section 4.3 (Annotation process). First, each annotator labels around 50 samples for each domain. The annotators are required to label the presence of hallucination, hallucination subtypes and locations, along with the corresponding explanations. For cases of inconsistent annotation, we invite experts to provide suggestions in a discussion. After that, annotators specify the application scope of each hallucination label as needed. Then the annotators take a vote for resolving the label-inconsistency of the first 50 samples in each domain. Second, the entire dataset is annotated according to this standard. The annotators label all the rest samples and vote for the inconsistent samples, following data checks by the experts. Third, we conduct data statistics of the whole dataset.
In the first step above, the discussion is organized in the form of online meetings. Annotators provide the inconsistent-labeled samples to experts (first 50 in each domain), after which all annotators and experts agree on a time for an online meeting discussion. Experts provide suggestions, and annotators modify the application scopes of hallucination labels based on the suggestions, thus making it more reliable.
The Price
The annotation time for each sample ranges from 2 to 10 (average 6.2) minutes. We pay each annotator 0.5 US dollars for annotating a sample and pay each expert 0.5 US dollars for checking a sample. This exceeds the local average hourly wage. Through the aforementioned approach, the quality of the annotations and the value of the benchmark are ensured. Thus, we consider it is greatly contributory to propose such a benchmark.
Label Consistency
@@ A.6 The Supplementary Details for Annotation
According to the above formulas, the calculated result for Fleiss’s Kappa of our benchmark is 0.8842, representing almost perfect agreement among all the annotators.
The Label Platform
Figure 12: The annotation interface on Label Studio.
A.7 The Application Scope of Hallucination Labels
Annotating the hallucination and its subtypes in this benchmark is a very challenging task. One of the reasons is that some hallucination subtypes in edge cases are defiant to differentiate. Therefore, in the first stage of annotation, we provide detailed definitions for each hallucination subtype. Below are the results of the discussion between the experts and the annotators.